OpenStudy (anonymous):

What is the point-slope equation of the line with slope -12 that goes through the point (5, 3)

OpenStudy (anonymous):

The point-slope formula is:\[y-y_1=m(x-x_1)\]Your point is (5,3) and your slope is m = -12, so\[y-3=-12(x-5)\] is the equation.